# Quasi-fuchsian flows and the coupled vortex equations

### v2

@@ 3. The bijection $\mathcal{G}$
But for the hyperbolic metric $h$, the weak stable bundle is given by $\ker(\beta_{1}+\psi_{1})$ and the claim follows; the constructed conjugacy is explicitly $df\circ\mathcal{I}:SM_{g}\to SM_{g_{1}}$. Running the same construction with $-A$ and $g_{2}$ (which amounts to switching $\lambda$ by $-\lambda$) we obtain a conjugacy between weak unstable foliations thus proving Items a, b, and c in Theorem 1.1. The proof of Item d is a direct consequence of [MP19, Theorem 5.5]. Finally, Item e is proved in Section 5 below. ∎
@@
The explicit map $\mathcal{I}$ provides an alternative and simpler way to compute the resonant 1-forms at zero for $F$; these were originally computed in [CP25, Theorem 1.5 and Section 8] using vertical Fourier analysis. Indeed, if we let $\text{Res}_{0}^{1}(g_{2})$ denote the space of resonant 1-forms of $g_{2}$ at zero in the kernel of contraction by the geodesic vector field (a well-known object), then the analogous space of resonant $1$-forms $\mathrm{Res}_{0}^{1}(F)$ corresponding to $F$ is given by pull-back by the map $\mathcal{I}$ corresponding to $-A$:
@@
It is easily computed that the Jacobian of $\mathcal{I}:W^{ws}(\zeta)\to W^{ws}(\tau)$ (where the two weak stable leaves are equipped with the orientations induced by the natural frames) is equal to $-K_{g}>0$, so $\mathcal{I}$ is orientation preserving. Hence we get that in fact $\tau^{\prime}$ and $\mathcal{I}(\zeta)$ are homotopic as curves with the same orientation. Since $\mathcal{I}(\zeta)$ has the same projection as $\zeta$ to $M$ (recall $\mathcal{I}$ covers the identity), we conclude that projection of $\tau^{\prime}$ belongs to the class $\mathfrak{c}$. Thus by uniqueness of closed orbits we get $\tau^{\prime}=\tau$, and the claim is proved.
@@
Observe that equations (1.6) and (1.7) imply that the thermostat flows associated with $(g,A)$ and $(g,-A)$ have the same marked length spectrum in the sense that the periods of closed orbits in a given free homotopy class match. More precisely, let $\mathfrak{c}$ be a free homotopy class on $M$, and let $\zeta_{\pm}$ be periodic orbits of the flow generating by $F_{\pm}:=X\pm\lambda V$ such that $\pi\circ\zeta_{\pm}$ belongs to the class $\mathfrak{c}$; here $\lambda=\operatorname{Im}(\pi_{2}^{*}A)$. Using that
@@
Since there is a Hölder orbit equivalence between the flows generated by $F_{\pm}$, a well-known application of the Livšic theorem gives that the flows are in fact Hölder conjugate (the conjugacy maps $\zeta_{+}$ to $\zeta_{-}$). This conjugacy cannot be upgraded to a smooth conjugacy as long as $A\neq 0$. Indeed, if we had a smooth conjugacy between the flows of $F_{\pm}$, then the weak stable foliations of the geodesic flows of $g_{1}$ and $g_{2}$ would be conjugate. In particular, the holonomies of the two foliations (i.e. the holonomy of foliated bundles, see [CC00, Chapter 2.1]) are conjugate, and these holonomies can be shown to agree with the representations of $\pi_{1}(M)\to\mathrm{PSL}(2,\mathbb{R})$ given by the respective hyperbolic structures. This implies that $[g_{1}]_{\mathrm{Teich}}=[g_{2}]_{\mathrm{Teich}}$ and thus $A=0$ by Theorem 1.1, Item d. The question of when a continuous conjugacy between dissipative Anosov flows on $3$-manifolds is smooth is analyzed in the recent preprint [GLH24]. Note that the Hölder conjugacy for quasi-Fuchsian flows must in fact swap SRB measures. This is a consequence of the observation that $r^{u}(g,A)=-r^{s}(g,-A)$ together with (5.1), the fact that $\zeta_{+}$ and $\zeta_{-}$ have the same period, and Bowen’s formula. This phenomenon is an obstruction to smooth conjugacy of transitive Anosov flows as explained in [GLH24].

### v3

@@ 3. The bijection $\mathcal{G}$
But for the hyperbolic metric $h$, the weak stable bundle is given by $\ker(\beta_{1}+\psi_{1})$ and the claim follows; the constructed conjugacy is explicitly $df\circ\mathcal{I}:SM_{g}\to SM_{g_{1}}$. Running the same construction with A-A and $g_{2}$ (which amounts to switching $\lambda$ by $\lambda$) we obtain a conjugacy between weak unstable foliations thus proving Items a, b, and c in Theorem 1.1. The proof of Item d is a direct consequence of [MP19, Theorem 5.5]. Finally, Item e is proved in Section 5 below. ∎
@@
The explicit map $\mathcal{I}$ provides an alternative and simpler way to compute the resonant 1-forms at zero for $F$; these were originally computed in [CP25, Theorem 1.5 and Section 8] using vertical Fourier analysis. Indeed, if we let $\text{Res}_{0}^{1}(g_{2})$ denote the space of resonant 1-forms of $g_{2}$ at zero in the kernel of contraction by the geodesic vector field (a well-known object), then the analogous space of resonant $1$-forms $\mathrm{Res}_{0}^{1}(F)$ corresponding to $F$ is given by pull-back by the map $\mathcal{I}$ corresponding to A-A:
@@
It is easily computed that the Jacobian of $\mathcal{I}:W^{ws}(\zeta)\to W^{ws}(\tau)$ (where the two weak stable leaves are equipped with the orientations induced by the natural frames) is equal to $K_{g}>0$, so $\mathcal{I}$ is orientation preserving. Hence we get that in fact $\tau^{\prime}$ and $\mathcal{I}(\zeta)$ are homotopic as curves with the same orientation. Since $\mathcal{I}(\zeta)$ has the same projection as $\zeta$ to $M$ (recall $\mathcal{I}$ covers the identity), we conclude that projection of $\tau^{\prime}$ belongs to the class $\mathfrak{c}$. Thus by uniqueness of closed orbits we get $\tau^{\prime}=\tau$, and the claim is proved.
@@
Observe that equations (1.6) and (1.7) imply that the thermostat flows associated with $(g,A)$ and (g,A)(g,-A) have the same marked length spectrum in the sense that the periods of closed orbits in a given free homotopy class match. More precisely, let $\mathfrak{c}$ be a free homotopy class on $M$, and let $\zeta_{\pm}$ be periodic orbits of the flow generating by $F_{\pm}:=X\pm\lambda V$ such that $\pi\circ\zeta_{\pm}$ belongs to the class $\mathfrak{c}$; here $\lambda=\operatorname{Im}(\pi_{2}^{*}A)$. Using that
@@
Since there is a Hölder orbit equivalence between the flows generated by $F_{\pm}$, a well-known application of the Livšic theorem gives that the flows are in fact Hölder conjugate (the conjugacy maps $\zeta_{+}$ to $\zeta_{-}$). This conjugacy cannot be upgraded to a smooth conjugacy as long as $A\neq 0$. Indeed, if we had a smooth conjugacy between the flows of $F_{\pm}$, then the weak stable foliations of the geodesic flows of $g_{1}$ and $g_{2}$ would be conjugate. In particular, the holonomies of the two foliations (i.e. the holonomy of foliated bundles, see [CC00, Chapter 2.1]) are conjugate, and these holonomies can be shown to agree with the representations of $\pi_{1}(M)\to\mathrm{PSL}(2,\mathbb{R})$ given by the respective hyperbolic structures. This implies that $g_{1}]_{\mathrm{Teich}}=[g_{2}]_{\mathrm{Teich}}$ and thus $A=0$ by Theorem 1.1, Item d. The question of when a continuous conjugacy between dissipative Anosov flows on $3$-manifolds is smooth is analyzed in the recent preprint [GLH24]. Note that the Hölder conjugacy for quasi-Fuchsian flows must in fact swap SRB measures. This is a consequence of the observation that $r^{u}(g,A)=-r^{s}(g,-A)$ together with (5.1), the fact that $\zeta_{+}$ and $\zeta_{-}$ have the same period, and Bowen’s formula. This phenomenon is an obstruction to smooth conjugacy of transitive Anosov flows as explained in [GLH24].